The annual budgeted conversion costs for a lean cell are $110,700 for 2,700 production hours. Each unit produced by the cell requires 18 minutes of cell process time. During the month, 1,170 units are manufactured in the cell. The estimated materials cost is $65 per unit. Provide the following journal entries:

a. Materials are purchased to produce 1,230 units.

b. Conversion costs are applied to 1,170 units of production.

c. 1,110 units are completed and placed into finished goods.
Answer #1

SOLUTION:

Particulars Debit Credit
Raw and In-Process invemtory 79,950
   Accounts payable 79,950
(1,230 units *$65)
Raw and In-Process inventory 14,391
Conversion Costs 14,391
[Conversion cost per hour = 110,700 / 2,700 = 41 per hour]
[Conversion cost applied = 1,170 * 18/60 * 41 = 14,391]
Finished Goods Inventory 13,718
   Raw and In-Process inventory 13,718
[Cost of units completed = 65 + 1,110*18/60*41 = 13,718]